OpenStudy (anonymous):

It takes 4 1/2 minutes for Elton John to listen to his favorite song on a record recorded at 45 r.p.m. (revolutions per minute). One day, Elton accidentally played this record at 33 1/3 r.p.m.'s. What was the number of revolutions required to play the song? What was the amount of time it took to play it at 33 1/3 r.p.m.?

OpenStudy (dumbcow):

revolutions = rate*time total revolutions = 4.5*45 = 202.5 time = revolutions/rate time = 202.5/33.33 = 6.075 min

OpenStudy (anonymous):

the total number of revolutions at 45 rpm = 45 * 9/2 = 202.5 revolutions. the number of revolutions will still be the same 33 1/3 rpms the time required would be 202.5*3/100 = 6.075 min.